# PAID: A Framework of Product-Centric Advertising Image Design

### v2

@@ Abstract.
Creating visually appealing advertising images is often a labor-intensive and time-consuming process. Is it possible to automatically generate such images using only basic product information—specifically, a product foreground image, taglines, and a target size? Existing methods mainly focus on parts of the problem and fail to provide a comprehensive solution. To address this gap, we propose a novel multistage framework called Product-Centric Advertising Image Design (PAID). It consists of four sequential stages to highlight product foregrounds and taglines while achieving overall image aesthetics: prompt generation, layout generation, background image generation, and graphics rendering. Different expert models are designed and trained for the first three stages: First, we use a visual language model (VLM) to generate background prompts that match the products. Next, a VLM-based layout generation model arranges the placement of product foregrounds, graphic elements (taglines and decorative underlays), and various nongraphic elements (objects from the background prompt). Following this, we train an SDXL-based image generation model that can simultaneously accept prompts, layouts, and foreground controls. To support the PAID framework, we create corresponding datasets with over 50,000 labeled images. Extensive experimental results and online A/B tests demonstrate that PAID can produce more visually appealing advertising images.
@@ 1. Introduction
Advertising images (as shown in Figure 1) are essential for commercial recommendation. Visually appealing and attractive images usually have high click-through rates (Wang et al., 2021) but are labor-intensive and time-consuming to create (Vashishtha et al., 2024). With the rapid advancement of generative methods (Menick and Kalchbrenner, 2018; Van Den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2021), it is now possible to automatically generate advertising images using only basic product inputs: a product foreground image, marketing taglines, and a target size.
@@
Currently, generating advertising images using only the basic inputs mentioned above has rarely been explored. Existing methods (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022) mainly focus on parts of the problem, such as background or layout generation, rather than offering a complete solution for integrating product foregrounds and taglines. Although these methods can be combined, they struggle to effectively highlight product images and taglines while maintaining visual appeal and harmony. Specifically, building on existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize predefined rules to create background prompts. These prompts are then fed into text-to-image inpainting models to complete backgrounds for product foregrounds. Layouts for taglines and decorative elements (collectively referred to as graphic elements) are predicted based on the images and taglines. Finally, attribute prediction and graphics rendering techniques are employed to add these graphic elements to the images. However, this process has several limitations: a) Lack of adaptability: It cannot generate background prompts that match the specific content and shape of the product, leading to poor foreground-background harmony. b) Fixed foreground positioning: The position of the foreground cannot be adjusted based on its characteristics and the background prompt, potentially resulting in inappropriate product size and poor composition. c) Restricted tagline placement: When adding taglines, the whole image content is already set, limiting the available space for tagline placement.
@@
Figure 2. Pipeline of PAID. It consists of four stages and generates advertising images centered around product information.
@@
In this paper, we propose a novel framework called Product-Centric Advertising Image Design (PAID) to enhance advertising image generation using only basic product inputs. While PAID remains a multistage process, we have refined its stages and task definitions to enhance foreground-background compatibility and the spatial arrangement of product foregrounds and taglines.
@@
As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create background prompts that better match the scene, placement, and angle based on product foregrounds. Following this, we propose a layout generation task that arranges the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements (objects to be generated from the background prompt). This arrangement is based on the product foreground, background prompt, marketing taglines, and target size. By arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the graphic layout, it becomes less constrained by the image content and allows for a more rational presentation of the product info. Furthermore, considering the nongraphic elements together helps minimize conflicts with taglines that could affect readability and attractiveness. Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2023) that can be controlled by input layouts. Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals. A LoRA (Hu et al., 2021) adaption strategy is adopted to balance conflicts and facilitate training. Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet (He et al., 2016), combined with a multi-scale training strategy to enhance training stability and efficiency. Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image. To build the framework and verify its effectiveness, we have collected and labeled over 50,000 images to establish datasets. In testing, PAID demonstrates its ability to produce advertising images with better visual quality. It has been deployed on two advertising recommendation scenarios of Taobao, and online A/B tests show that its generated images are more appealing to users.
@@
  • We propose a novel framework PAID for advertising image design given only a product foreground image, taglines, and a target size. It centers on the product foreground and taglines to enhance the overall product prominence and improve the image aesthetics.
@@ 2.1. Advertising Image Design
In recent years, many methods have been proposed to automate advertising image design (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a). They can be divided into two subcategories: background-based and foreground-based methods. Background-based methods predict graphic layouts based on complete images (not only product foregrounds), followed by attribute prediction and graphics rendering modules to create the final advertising image. CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage image info and generate layouts. RADM (Li et al., 2023b) is the first to simultaneously consider the image and tagline content to derive layouts with a diffusion model. PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout. AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a complete image, and then predicts the attributes of graphic elements and renders them onto the image. For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, including the number and length of taglines.
In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines. It subsequently designs a module network to outpaint the background and render the taglines according to the given background prompt and generated layout. The tagline content is not limited, and the layout is more flexible. However, P&R ignores the harmony between the background prompt and the generated layout. The layout may conflict with background generation. In our PAID framework, the overall layout of graphic elements, product foreground, and other nongraphic elements is generated jointly according to the product foreground, background prompt, and tagline contents, aiming to achieve the best harmony among them.
@@ 2.2. Controllable Image Generation
There have been significant developments in text-to-image models (Rombach et al., 2021; Podell et al., 2023). To control instances and further improve image quality, some methods introduce layout guidance, dividing into training-free and training-based categories. Training-free methods manipulate the attention map (Kim et al., 2023) or compute attention-based loss during de-noising (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024; Bar-Tal et al., 2023). Training-based methods, like GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a), modify network architecture, incorporating layout data by adding gated self-attention layers.
For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized. The inpainting technique is employed to preserve product characteristics, while the prompt is used to describe the background. The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2021) or employing an inpainting ControlNet (Zhang et al., 2023). To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control. Currently, there is limited exploration in this area. SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2021) that can accept layout and foreground controls. However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images. Therefore, we use several strategies when training the layout-controlled inpainting model to ensure training stability, generation quality, and efficiency.
@@ 3. PITA Dataset and PIL Dataset
For automatic image design of product-centric advertising, we collect a large-scale Product-Centric Image-Tagline Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022). 1,000 of them are used for testing. PITA is a dataset of varying sizes, featuring images with four distinct aspect ratios: 0.684, 1.0, 0.667, and 0.75. Each image is labeled with prompts (foreground caption and background caption), a product mask, and a layout of graphic and nongraphic elements. Each element is represented with a type and a bounding box (bbox). The graphic elements contain “Logo”, “Tagline”, and “underlay”. Some examples of data visualization are in the appendix.
For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset. The images do not contain graphic elements and are labeled with prompts, product masks, and layouts of nongraphic elements. PIL contains 12,247 samples in total and 1,000 of them are divided for test.
@@ 4.1. Overall Framework
As illustrated in Figure 2, our proposed PAID framework consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
@@ 4.2. Prompt generation
Input data format. For the input of LLM, we first describe the prompt generation task and ask for foreground and background descriptions. For the image input, we crop out the product foreground and feed it into the visual encoder.
Output data format. To enable downstream layout and background image generation tasks, and to better understand the relationship between the foreground and background, we use the prompt generation model $\pi$ to simultaneously predict the foreground description $p_{fore}$ and the background prediction $p_{back}$. We format the output in JSON as it is compatible with the pre-trained VLM and simplifies the following analysis. The output JSON is structured as $p=(p_{fore},p_{back})$.
Training scheme. A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2024a). During pre-training, the projector learns to align features between visual and text modalities. In the SFT stage, the projector and LLM are further trained for visual understanding tasks. Since we utilize the VLM trained after the SFT stage, the model can understand input images and follow instructions. We only fine-tune the LLM branch with our dataset to adapt the model to the prompt generation task, using cross-entropy loss as the objective function. The model $\pi_{prompt}$ takes the foreground image $I_{fore}$ and predefined instructions as input, performing two main tasks: generating a foreground description $p_{fore}$ and predicting the background description $p_{back}$. The prompt generation process is as follows:
(1) $p=\pi_{prompt}(I_{fore}).$
@@ 4.3. Layout generation
Figure 3. The framework of our prompt and layout generation model and data format.
Input data format. For the input to the LLM, we describe the layout generation task and organize input conditions in JSON format. The image input remains the same as in prompt generation. A target canvas aspect ratio is given as a condition to support multi-scale layout design. The aspect ratio of the product foreground is provided to help keep its shape. Foreground and background prompts are also given for generating the image layout. For the pre-defined taglines, we list them with empty locations and expect the model to complete them. If a logo is needed, we include it as a condition in the graphic layout section and specify its aspect ratio.
Except for these basic conditions, we also take product characteristics into account. In E-commerce scenarios, some products should not be occluded for better display. While for others, like clothing or human-hold items, partial occlusion is acceptable. To differentiate between these, we introduce a Class-Conditioned Layout Prediction (CCLP) strategy. CCLP provides clear information about the product class and whether foreground occlusion is allowed. For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1]. The bounding boxes of taglines should never occlude the subject”. For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2]. The bounding boxes of taglines are allowed to occlude the subject”.
@@
Output data format. The output layout is structured in JSON format with three parts: subject layout, other nongraphic layout, and graphic layout. To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR). It is expressed as [x, y, h*r1/r2, h]. (x,y) is the normalized center point. w and h are the normalized width and height. r1 and r2 are the actual aspect ratio of the foreground and canvas, respectively. Each value ranges from 0 to 1 with three decimal places. The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox). The graphic layout defines each element by type, content, and bounding box, covering tagline, underlay, and logo. The content key is used only for tagline types to specify its semantic content. Bounding boxes are listed as [x, y, w, h]. The model is expected to supplement underlay adaptively. If logos are required, it is represented in the same way as the product for ratio keeping.
Training scheme. Same as the prompt generation module, we only fine-tune the LLM branch with the visual encoder and projector frozen. The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates. Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and prevent the model from memorizing the input sequence.
@@ 4.4. Background Image generation
Layout control. Suppose that there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$. We map these into $N$ embeddings as layout input. Here, $n_{i}$ is the element name, and $l_{i}$ is its location. Concretely, we use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$. $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$. Next, $\textbf{e}_{n,i}$ and $\textbf{e}_{l,i}$ are concatenated and fed into a trainable MLP layer to represent the layout information of the $i$-th element, denoted as $\textbf{e}_{i}$. $\textbf{e}=\{\textbf{e}_{i}|i=0,1,...,N-1\}$ is the layout embeddings. A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings. Supposing there are $M$ visual embeddings (V) after the self-attention layer. The output of the GSA layer is
@@
where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
Considering that the layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet. We refer to this strategy as Deep Layer Control (DLC). Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks. This reduces the parameters and inference latency, and has no negative impact on generation quality. Additionally, we adopt a LoRA Adaptation Training (LAT) strategy in which trainable LoRA (Hu et al., 2021) layers are added to UNet, to bridge the gap between the base model and added GSA layers. This further accelerates convergence and increases generation quality. Figure 4 describes how layout control is achieved.

### v3

# T-Stars-Poster: A Framework for Product-Centric Advertising Image Design

@@ Abstract.
Creating advertising images is often a labor-intensive and time-consuming process. Can we automatically generate such images using basic product information like a product foreground image, taglines, and a target size? Existing methods mainly focus on parts of the problem and lack a comprehensive solution. To bridge this gap, we propose a novel product-centric framework for advertising image design called T-Stars-Poster. It consists of four sequential stages to highlight product foregrounds and taglines while achieving overall image aesthetics: prompt generation, layout generation, background image generation, and graphics rendering. Different expert models are designed and trained for the first three stages: First, a visual language model (VLM) generates background prompts that match the products. Next, a VLM-based layout generation model arranges the placement of product foregrounds, graphic elements (taglines and decorative underlays), and various nongraphic elements (objects from the background prompt). Following this, an SDXL-based model can simultaneously accept prompts, layouts, and foreground controls to generate images. To support T-Stars-Poster, we create two corresponding datasets [^1] The primary dataset is available at https://github.com/alimama-creative/PITA-dataset. with over 50,000 labeled images. Extensive experiments and online A/B tests demonstrate that T-Stars-Poster can produce more visually appealing advertising images.
@@ 1. Introduction
Advertising images (as shown in Figure 1) are essential for commercial recommendation. Visually appealing images usually have high click-through rates (Wang et al., 2021) but are labor-intensive and time-consuming to create (Vashishtha et al., 2024). With the rapid advancement of generative methods (Menick and Kalchbrenner, 2018; Van Den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2021), it is now possible to automatically generate advertising images using only basic product inputs: a product foreground image, marketing taglines, and a target size.
@@
Currently, generating advertising images using only the basic inputs has rarely been explored. Existing methods (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022) mainly focus on parts of the problem, such as background or layout generation, rather than offering a complete solution for integrating product foregrounds and taglines. Although these methods can be combined, they struggle to effectively highlight product images and taglines while maintaining visual appeal and harmony. Specifically, building on existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize predefined rules to create background prompts, which are then fed into text-to-image inpainting models to complete backgrounds for product foregrounds. Layouts for taglines and decorative elements (collectively referred to as graphic elements) are predicted based on the images and taglines. Finally, attribute prediction and graphics rendering techniques are employed to add these graphic elements to the images. However, this process has several limitations: a) Lack of adaptability: The background prompts cannot match the specific content and shape of the product, leading to poor foreground-background harmony. b) Fixed foreground positioning: The foreground position cannot be adjusted based on the product characteristics and the background prompt, potentially resulting in inappropriate product size and poor composition. c) Restricted tagline placement: The whole image content is set before taglines are added, limiting space for tagline placement.
@@
Figure 2. Pipeline of T-Stars-Poster. It consists of four stages and generates advertising images centered around product information.
@@
In this paper, we propose a novel product-centric framework called T-Stars-Poster for advertising image generation using only basic product inputs. We refine the stages and task definitions to enhance foreground-background compatibility and the spatial arrangement of product foregrounds and taglines.
@@
As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to create background prompts that match the scene, placement, and angle with product foregrounds. Next, a layout generation task arranges the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements (objects to be generated from the background prompt). This arrangement is based on the product foreground, background prompt, taglines, and target size. By arranging the product foreground and taglines simultaneously, rather than generating the whole image before deciding the graphic layout, we reduce layout constraints and present product information more effectively. Furthermore, considering nongraphic elements together helps minimize conflicts with taglines that could affect readability and attractiveness. We then train an inpainting text-to-image model based on SDXL (Podell et al., 2023) that can be guided by input layouts. Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals. A LoRA (Hu et al., 2021) adaption strategy is adopted to balance conflicts and facilitate training. Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet (He et al., 2016), along with a multi-scale training strategy to enhance stability and efficiency. Finally, we use an attribute prediction and graphics rendering module to overlay graphic elements onto the image. To build the framework and verify its effectiveness, we have collected and labeled over 50,000 images for datasets. In test, T-Stars-Poster shows it can produce advertising images with better visual quality. It has been deployed on two advertising recommendation scenarios of Taobao, and online A/B tests show that its generated images are more appealing to users.
@@
  • We propose a framework T-Stars-Poster for advertising image design given only a product foreground, taglines, and a target size. It centers on product foregrounds and taglines to improve the product prominence and image aesthetics.
@@ 2.1. Advertising Image Design
Many methods have been proposed to automate advertising image design (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a), falling into two categories: background-based and foreground-based. Background-based methods (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023) predict graphic layouts for complete images (not product foregrounds), followed by attribute prediction and graphics rendering to create the final image. These methods face spatial constraints due to fixed image content, limiting tagline content. In contrast, P&R (Li et al., 2023a) is a foreground-based method using both foreground images and tagline content to predict layout, then outpainting the background and rendering taglines according to the given background prompt and generated layout. The tagline content and layout are more flexible. However, P&R ignores the harmony between the background prompt and the generated layout. The layout may conflict with background generation. Our framework generates the overall layout for graphic elements, product foreground, and other elements jointly, using product foregrounds, background prompts, and tagline content to ensure harmony.
@@ 2.2. Controllable Image Generation
There have been significant developments in text-to-image models (Rombach et al., 2021; Podell et al., 2023). To control instances and improve image quality, some methods introduce layout guidance, dividing into training-free and training-based categories. Training-free methods adjust the attention map or compute attention-based loss during inference (Kim et al., 2023; Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024; Bar-Tal et al., 2023). Training-based methods (Li et al., 2023c; Wang et al., 2024a) add gated self-attention layers to models and train them with corresponding data.
For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models with inpainting capabilities are often utilized. Inpainting preserves foregrounds and prompts describe the backgrounds. The foreground image is integrated using extra channels (Rombach et al., 2021) or an inpainting ControlNet (Zhang et al., 2023). To better manage the layout and enhance image quality, we combine inpainting with layout control. Currently, there is limited exploration in this area. SceneBooth (Chai et al., 2025) is the first to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2021). However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images. Therefore, we use several strategies to ensure training stability, generation quality, and efficiency.
@@ 3. PITA Dataset and PIL Dataset
We collect a Product-Centric Image-Tagline Advertising (PITA) dataset with 38,017 samples from Taobao and the CGL dataset (Zhou et al., 2022). 1,000 of them are for test. PITA covers major categories of Taobao, such as clothing, food, furniture, and electronics. It features images with four distinct aspect ratios: 0.684, 1.0, 0.667, and 0.75. In the collection process, we exclude images with messy backgrounds, plain colors without shadows, unattractive stickers, close-ups, poorly presented taglines, or cluttered elements. Each has labels for prompts (foreground and background captions), a product mask, and a layout. Each element is represented with a type and a bounding box (bbox). Graphic elements contain “Logo”, “Tagline”, and “Underlay”. Initial annotations are generated using automatic methods such as DAMO Academy’s matting API (Academy, 2024) for foreground extraction, GPT-4o (OpenAI, 2024) for image captioning, a detection model (Dai et al., 2021) trained with the CGL dataset and Grounding-DINO (Liu et al., 2024c) for element detection, and an OCR model (PaddlePaddle, 2023) for tagline extraction. Annotations are then manually reviewed and corrected for accuracy.
We also develop a Product-Centric Image Layout (PIL) dataset for layout-controlled background inpainting. It contains images without graphic elements and includes 12,247 samples, with 1,000 reserved for testing. The filtering and annotation process is the same as the PITA dataset, excluding graphic elements.
@@ 4.1. Overall Framework
As illustrated in Figure 2, our proposed T-Stars-Poster framework consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
@@ 4.2. Prompt Generation
Input data format. We describe the prompt generation task and request foreground/background descriptions. Meanwhile, we feed the product foreground into the visual encoder.
Output data format. To support layout and background image generation tasks and understand the foreground-background relationship, the prompt generation model $\pi$ predicts both the foreground description $p_{fore}$ and background description $p_{back}$ simultaneously. We format the output in JSON as it is compatible with the pre-trained VLM and simplifies the following analysis. The output JSON is structured as $p=(p_{fore},p_{back})$.
Training scheme. We use a VLM trained after supervised fine-tuning (SFT) (Liu et al., 2024a), allowing the model to understand images and follow instructions. We only fine-tune the LLM part with our dataset for the prompt generation task, using cross-entropy loss. The model $\pi_{prompt}$ takes a foreground image $I_{fore}$ and predefined instructions to generate a foreground description and predict a background description. The prompt generation process is as follows:
(1) $p=(p_{fore},p_{back})=\pi_{prompt}(I_{fore}).$
@@ 4.3. Layout Generation
Figure 3. The framework of prompt and layout generation.
Input data format. For LLM input, we describe the layout generation task and organize conditions in JSON format. The image input is the same as in prompt generation. A target canvas aspect ratio is provided to support multi-scale design. The aspect ratio of the product foreground is given to keep its shape. Foreground/background prompts are included for image layout generation, and taglines are listed with empty locations for the model to fill in. If a logo is needed, it will be included with its aspect ratio.
We also take product characteristics into account. In e-commerce, some products need to be fully visible, while partial occlusion is acceptable for others, like clothing. To address this, we use a Class-Conditioned Layout Prediction (CCLP) strategy that indicates the product class and whether occlusion is allowed. For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1]. The bounding boxes of taglines should never occlude the subject”. For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2]. The bounding boxes of taglines are allowed to occlude the subject”.
@@
Output data format. The output layout is structured in JSON format with three parts: subject layout, other nongraphic layout, and graphic layout. To maintain the foreground shape, we propose a Ratio-Keeping Bbox Representation (RKBR): [x, y, h*r1/r2, h]. (x,y) is the normalized center point. w and h are the normalized width and height. r1 and r2 are the actual aspect ratio of the foreground and canvas, respectively. Each value ranges from 0 to 1. The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox). The graphic layout defines each element by type, content, and bounding box, covering the tagline, underlay, and logo. The content key is used only for tagline types to specify its semantic content. Bounding boxes are listed as [x, y, w, h]. The model is expected to supplement underlay adaptively. If logos are required, it is represented in the same way as the foreground for ratio keeping.
Training scheme. Like the prompt generation module, we fine-tune only the LLM branch. The output JSON predicts layouts in a raster scan order, following ascending top-left coordinates. We also shuffle taglines in the input JSON for data augmentation to prevent the model from memorizing the sequence.
@@ 4.4. Background Image Generation
Layout control. Suppose an image has $N$ elements, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$. We map these into $N$ embeddings as layout input. Here, $n_{i}$ is the element name, and $l_{i}$ is its location. We use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$. $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$. Next, $\textbf{e}_{n,i}$ and $\textbf{e}_{l,i}$ are concatenated and fed into a trainable MLP layer to represent the layout information of the $i$-th element, denoted as $\textbf{e}_{i}$. $\textbf{e}=\{\textbf{e}_{i}|i=0,1,...,N-1\}$ is the layout embeddings. A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings. If there are $M$ visual embeddings (V) after the self-attention layer, the GSA output is
@@
where $\gamma$ is a learnable parameter, catcat means concatenation, and $SA$ means self-attention.
Since layout relates to semantic-level information and not high-frequency details, GSA layers are only added in the deep blocks of UNet. We refer to this strategy as Deep Layer Control (DLC). Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks. Additionally, we use a LoRA Adaptation Training (LAT) strategy, adding LoRA (Hu et al., 2021) layers to UNet to bridge the gap between the base model and added GSA layers. These further accelerate convergence and increase generation quality. Figure 4 describes how layout control is achieved.